# Can Open-source LLMs Enhance Data Augmentation for Toxic Detection?: An Experimental Study

### v1

@@ Abstract
High-quality, diverse harmful data is essential to addressing real-time applications in content moderation. Current state-of-the-art approaches to toxic content detection using GPT series models are costly and lack explainability. This paper investigates the use of prompt engineering and fine-tuning techniques on open-source LLMs to enhance harmful data augmentation specifically for toxic content detection. We conduct a two-stage empirical study, with stage 1 evaluating six open-source LLMs across multiple datasets using only prompt engineering and stage 2 focusing on fine-tuning. Our findings indicate that Mistral can excel in generating harmful data with minimal hallucination. While fine-tuning these models improves data quality and diversity, challenges such as data duplication and overfitting persist. Our experimental results highlight scalable, cost-effective strategies for enhancing toxic content detection systems. These findings not only demonstrate the potential of open-source LLMs in creating robust content moderation tools. The application of this method in real industrial scenarios further proves the feasibility and efficiency of the fine-tuned open-source LLMs for data augmentation. We hope our study will aid in understanding the capabilities and limitations of current models in toxic content detection and drive further advancements in this field.
@@ 1 Introduction
With the development of large language models Brown et al. (2020), there is a significant amount of work currently using LLMs for data augmentation Whitehouse et al. (2023); Ye et al. (2022); Wang et al. (2023). This study investigates the application of prompt engineering and fine-tuning techniques on open-source large language models to enhance the augmentation of harmful data, crucial for improving toxic content detection systems. Data augmentation for toxic detection is vital because real-world datasets often lack the diversity and volume needed to train robust models Garg et al. (2023). Current challenges in toxic detection include the need for real-time processing Oikawa et al. (2022), where toxic comments must be filtered out swiftly to prevent user exposure. Existing methods using LLMs to directly detect toxic content, while effective, often suffer from slow inference speeds and high computational resource requirements, making them unsuitable for systems without GPU resources Kumarage et al. (2024). Additionally, methods based on the GPT Brown et al. (2020) series, despite their good performance, face issues of high cost and lack of explainability Zhao et al. (2024). Content safety classifiers also often face urgency and specificity across various scenarios due to differing definitions of sensitive topics among users, and obtaining harmful training data is inherently challenging. As a result, many products require a standalone data augmentation module, which can be used as an independent service to meet the training data generation needs of various applications. Using GPT models is both costly and slow Hui et al. (2024), whereas open-source LLMs can make data generation faster and cheaper with comparable effectiveness Ding et al. (2024). This work aims to address these challenges by optimizing the use of open-source LLMs for generating high-quality, diverse, and compliant harmful data, thus enhancing the reliability and accuracy of toxic content detection systems. This flexibility helps enterprises respond to emergency topics and cater to user-specific customization needs. Additionally, the fine-tuned models have already been deployed in our API, powering numerous use cases such as chatbot content safety, live chat, and children’s custom content channels.
Our experimental approach includes a comprehensive evaluation of various open-source LLMs, including Mistral Jiang et al. (2023), LLaMa2 Touvron et al. (2023), Vicuna Chiang et al. (2023), Falcon Almazrouei et al. (2023), Bloom Workshop et al. (2023), and Gemma Team et al. (2024), across different prompt engineering strategies Schulhoff et al. (2024); Li (2023) and fine-tuning configurations. The research is divided into two stages. In the first stage, we focus on prompt engineering by comparing LLMs, designing effective prompts, and experimenting with few-shot techniques to balance quality and diversity in the generated data Ding et al. (2024). We have found that relying solely on prompt engineering often results in poor-quality generation due to the model’s internal safety alignments. Therefore, we conducted a second stage of experiments using data for fine-tuning.
In the second stage, we explore fine-tuning LLMs using internal datasets and various configurations, such as epoch settings and data mixing experiments. Results show that fine-tuning can mitigate the generation of hallucination by LLMs and improve performance in generating harmful data. However, issues like data duplication and overfitting are observed, suggesting that a careful balance in training epochs and data variety is essential. The fine-tuned versions are already being used in our product.
Our findings provide critical insights into optimizing open-source LLMs for harmful data generation, thereby enhancing the robustness and accuracy of toxic content detection systems. By addressing key challenges such as data quality, diversity, and compliance, our work offers scalable, cost-effective strategies. These strategies are essential for improving the safety and reliability of automated content moderation tools, making them more effective in real-world applications. This research not only informs the development of a larger harmful detection system but also validates the efficiency of fine-tuned open-source LLMs in industrial scenarios. Our main contributions are summarized as follows:
  • Our study is among the first to explore prompt engineering and fine-tuning techniques on open-source LLMs to enhance harmful data augmentation for toxic content detection.
  • Our findings highlight the potential of fine-tuned open-source LLMs to address real-time processing challenges, offering cost-effective, scalable strategies for improving automated content moderation tools and informing the future direction of larger harmful detection systems.
  • Fine-tuned models have been deployed for our product, powering various use cases such as chatbot content safety, live chat, and children’s custom content channels and limitations of the existing LLMs are analyzed and discussed using real-world data.
@@ 2 Methodology
Our approach leverages prompt engineering and fine-tuning techniques on open-source LLMs for harmful data augmentation, enhancing toxic content detection systems. The methodology consists of two main stages: prompt engineering and fine-tuning.
@@ 2.1 Prompt Engineering
where Role defines the context, Requirement specifies the task, and Few-Shot Examples provide examples to guide the model. In the initial phase of our study, we focused solely on prompt engineering. This involved creating various prompt templates tailored to elicit specific types of harmful data. Each prompt was carefully crafted following work by Zhou et al. (2023) to ensure clarity and relevance to the desired output. Despite the meticulous design of these prompts, we encountered significant challenges. The primary issue was the model’s internal safety alignments, which are designed to prevent the generation of harmful content. While these alignments are crucial for ethical AI usage, they posed a hurdle for our specific goal of generating harmful data for augmentation purposes. Consequently, the generated data often lacked the quality and diversity needed for effective toxic content detection. To overcome these challenges, we experimented with various prompt configurations, including adjusting the few-shot examples and altering the specificity of the requirements. However, these adjustments had limited success in mitigating the internal safety mechanisms. After testing, the final prompt template are showed in A. Based on these findings, we concluded that relying solely on prompt engineering was insufficient for our needs. This realization led us to incorporate a second stage involving fine-tuning, where we aimed to enhance the model’s capability to generate high-quality harmful data by updating its weights using carefully curated training datasets.
@@ 2.2 Fine-Tuning
Based on the results from prompt engineering, we believed that fine-tuning would provide a better solution. We used LoRA methods Hu et al. (2022) for fine-tuning to update the weights of $\mathcal{M}$ using a training dataset $\mathcal{D}_{\text{train}}$ to improve the quality and diversity of $\mathcal{D}_{\text{harmful}}$. The fine-tuning objective $\mathcal{L}$ minimizes the cross-entropy loss:
$\mathcal{L}=-\sum_{i=1}^{N}\left(y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i})\right)$
where $y_{i}$ is the true label, $\hat{y}_{i}$ is the predicted probability, and $N$ is the number of samples. After fine-tuning, Mistral performed the best compare with baselines; the detailed experiment result are shown in 4. The fine-tuned versions are already being used internally.
@@ 2.3 Algorithm
The overall algorithm for harmful data augmentation is outlined in Algorithm 1.
Algorithm 1 Harmful Data Augmentation using Open-Source LLMs
1:Open-source LLM $\mathcal{M}$, training dataset $\mathcal{D}_{\text{train}}$, prompt templates $\mathcal{P}$
2:Augmented harmful data $\mathcal{D}_{\text{augmented}}$
@@
5:Fine-tune LLM $\mathcal{M}$ on $\mathcal{D}_{\text{train}}$ using LoRA methods
@@
7:     Generate harmful data samples $\mathcal{D}_{\text{harmful},i}$ using the fine-tuned LLM $\mathcal{M}$ and $\mathcal{P}_{i}$
@@
9:Combine all generated samples $\mathcal{D}_{\text{harmful}}=\bigcup_{i}\mathcal{D}_{\text{harmful},i}$
10:for each batch $B_{i}$ do
11:     Generate augmented data $\mathcal{D}_{\text{batch},i}$ using fine-tuned model $\mathcal{M}$
12:     Add $\mathcal{D}_{\text{batch},i}$ to $\mathcal{D}_{\text{augmented}}$
@@
14:return $\mathcal{D}_{\text{augmented}}$
Where $\mathcal{D}_{\text{train}}$ is the total training data for finetuning, $\mathcal{D}_{\text{harmful}}$ is the data samples (output) generated by the finetuned LLM, and $\mathcal{D}_{\text{augmented}}$ is the final dataset for downstream models training. Note that $\mathcal{D}_{\text{augmented}}$ include both $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{harmful}}$.
@@ 2.4 Evaluation
We evaluate the performance of the augmented data using F1-score and Accuracy metrics on a smaller MLP classifier(more detailed in 3.2 section).
@@ 3.1 Datasets
The following proprietary datasets are used for training models to identify various types of harmful content. Each dataset contains binary classification labels, with texts labeled as either positive or negative (non-harmful). The labels are based on multi-round human evaluation. The sources of the datasets include online public datasets, company services collection, and other sources. Detailed data numbers are listed in Table 1. We also prepare a 2500 entries evaluation dataset from each dataset.
@@
Table 1: Detailed Dataset Numbers
@@ 3.2 Models
We conducted a two-stage test to evaluate and fine-tune different language models for identifying harmful content.
@@
In the first stage, we employed prompt engineering and tested six different models: Mistral Jiang et al. (2023), LLaMa2 Touvron et al. (2023), Vicuna Chiang et al. (2023), Falcon Almazrouei et al. (2023), Bloom Workshop et al. (2023), Team et al. (2024). Stage 1 result can be found in 4. After evaluating the performance of these models, we determined that Mistral and Vicuna performed best in the prompt engineering stage. Therefore, we selected these two models for the fine-tuning stage.
@@
In the second stage, we fine-tuned the Mistral and Vicuna models using the LoRA method Hu et al. (2022). After comparing latency and cost we decided to use a smaller model Mixtral for further exploration.
@@ 3.3 Implementation
The fine-tuning process was conducted on 4*NVIDIA A100 GPUs Azure cluster. We utilized the Huggingface Transformers library for our fine-tuning experiments. The specific model architecture and pre-trained weights used in our study are available at Huggingface, as referenced in Wolf et al. (2020). The fine-tuning stage specified was Supervised Fine-Tuning (SFT). Adapter is a specific Low-Rank Adaptation Hu et al. (2022) fine-tuning setup. Both the maximum number of new tokens generated and the cutoff length for the input sequences were set to 4096. The evaluation batch size per device was set to 8. The results of the fine-tuning process, including detailed performance metrics and qualitative analyses, can be found in Section 4.
@@
For the downstream model evaluation, we utilized a Multilayer Perceptron Gardner and Dorling (1998) classifier with two hidden layers, each consisting of 600 neurons. The primary reason for selecting an MLP is due to its applicability in scenarios where GPU resources are limited. In actual applications, many implementations operate without the support of GPUs. Therefore, in a GPU-constrained environment, an MLP serves as a practical and efficient choice for classification tasks.
@@ 4 Result
The results of our study demonstrate the efficacy of both prompt engineering and fine-tuning strategies in enhancing the performance of open-source LLMs for harmful data augmentation. We evaluated different models using a two-stage process: prompt engineering and fine-tuning.
@@ 4.1 Stage 1: Prompt Engineering
The first stage of our evaluation focused on assessing the performance of various Open Source LLMs for data augmentation using prompt engineering. The primary goal was to explore the quality and effectiveness of existing LLMs in generating augmented data for political content. Table 2 presents the performance comparison of the six different models evaluated in Stage 1. The success rates for political and hate content are reported, along with the results of human evaluation.
@@
Hate Mistral 0.403 0.621
GPT4(baseline) 0.401 0.620
@@
hate_epoch5 0.755 0.788
hate_epoch3 0.671 0.740
hate_epoch2 0.653 0.731
hate_epoch1 0.587 0.702
@@ 4.2 Stage 2: Fine-tuning
In the second stage, we first explored fine-tuning the models using only one dataset. The results of this process are shown in Table 4. Inspired by Hui et al. (2024) ,Yu et al. (2024) and Qi et al. (2023), we then implemented a data mixing strategy, leveraging multiple datasets for fine-tuning. The results of fine-tuning the models on five different datasets are presented in Table 3. We also performed an ablation study on different design choices which are listed in Appendix C. The "Mixture" versions refer to fine-tuning data that is a mix of hate, violence, and sex datasets. The "Mix_GPT" versions refer to a combination of the mixed data with GPT-generated positive samples. For all fine-tuning processes, we used 3000 positive and 3000 negative samples, following the Alpaca instruction template.
Our results show that fine-tuning significantly improves the performance of LLMs in generating harmful data. The data mixing strategy helps in mitigating issues like data duplication and overfitting. The fine-tuned models exhibited higher F1 scores and accuracy across different datasets, demonstrating their robustness and effectiveness in real-world applications. One surprising finding was the substantial improvement in model performance when using the mixed data strategy. This suggests that diverse training data can help models generalize better and produce more accurate and reliable outputs. The Mixture models with Epoch 3 and Epoch 5 showed particularly notable improvements, indicating that sufficient training epochs are critical for achieving optimal performance.

### v2

# Can Open-source LLMs Enhance Data Synthesis for Toxic Detection?: An Experimental Study

@@ Abstract
Effective toxic content detection relies heavily on high-quality and diverse data, which serves as the foundation for robust content moderation models. This study explores the potential of open-source LLMs for harmful data synthesis, utilizing prompt engineering and fine-tuning techniques to enhance data quality and diversity. In a two-stage evaluation, we first examine the capabilities of six open-source LLMs in generating harmful data across multiple datasets using prompt engineering. In the second stage, we fine-tune these models to improve data generation while addressing challenges such as hallucination, data duplication, and overfitting. Our findings reveal that Mistral excels in generating high-quality and diverse harmful data with minimal hallucination. Furthermore, fine-tuning enhances data quality, offering scalable and cost-effective solutions for augmenting datasets for specific toxic content detection tasks. These results emphasize the significance of data synthesis in building robust, standalone detection models and highlight the potential of open-source LLMs to advance smaller downstream content moderation systems. We implemented this approach in real-world industrial settings, demonstrating the feasibility and efficiency of fine-tuned open-source LLMs for harmful data synthesis.
@@ 1 Introduction
In toxic content detection, (1) labeling real-world datasets is both expensive and labor-intensive, with the substantial volume of data required further compounding the challenge—our costs have reached millions of dollars. (2) Ensuring dataset diversity and quality remains a persistent issue, often hindering the development of robust detection models. (3) The inherent ambiguity and subjectivity of the term "harmful"—which varies across topics and annotator perspectives—adds another layer of complexity. Data synthesis presents a promising approach to mitigating these challenges. Moreover, many product applications demand real-time toxic content detection with stringent latency requirements (e.g., latency $\leq$ 0.01 seconds). Directly using SOTA Podolak et al. (2024); Fillies and Paschke (2024) LLM-based approaches might not reliably meet these constraints, making LLM-generated data synthesis for training smaller, more robust classification models a practical solution.
This study explores the use of prompt engineering and fine-tuning techniques with open-source LLMs to improve harmful data synthesis. Existing methods, such as ToxiGen Hartvigsen et al. (2022) and Toxicraft Hui et al. (2024) provide useful frameworks but rely on GPT-based frameworks for data generation, which face issues of cost and scalability. For example, ToxiGen employs GPT3 within the ALICE framework, producing high-quality datasets but at a high computational expense. These limitations highlight the need for alternative solutions. Our study addresses these challenges by investigating how open-source LLMs can facilitate harmful data synthesis. Open-source models offer comparable effectiveness to proprietary systems while reducing computational costs and enhancing explainability Ding et al. (2024). Many products can benefit from standalone data synthesis modules that generate training data for domain-specific and time-sensitive content safety applications, making open-source LLMs a scalable and cost-effective choice.
Our experimental approach includes a comprehensive evaluation of six open-source LLMs: Mistral Jiang et al. (2023), LLaMa2 Touvron et al. (2023), Vicuna Chiang et al. (2023), Falcon Almazrouei et al. (2023), Bloom Workshop et al. (2023), and Gemma Team et al. (2024), explored across diverse prompt engineering strategies Schulhoff et al. (2024); Li (2023) and fine-tuning configurations. The research is divided into two key stages: In the first stage, we focus on prompt engineering by comparing LLMs, designing effective prompts, and experimenting with few-shot techniques to balance quality and diversity in the generated data Ding et al. (2024). However, we observed that relying solely on prompt engineering often results in suboptimal outputs due to the model’s inherent safety alignments, which limit harmful data generation. Consequently, we conducted a second stage of experiments involving fine-tuning. In the second stage, we fine-tuned the LLMs using proprietary datasets and explored various configurations, such as epoch settings and data mixing strategies. Our results demonstrate that fine-tuning effectively mitigates issues like hallucination in LLMs and enhances the quality of harmful data generation. However, challenges such as data duplication and overfitting were identified, emphasizing the need for a careful balance between training epochs and dataset diversity.
In this study, we do not aim to develop a new state-of-the-art model in toxic content detection but rather to explore and experiment with intuitive, effective ideas tailored for industrial applications of toxic content moderation systems. These efforts provide key insights into optimizing open-source LLMs for harmful data synthesis, enhancing both the robustness and accuracy of toxic content detection systems. Additionally, this study validates the practicality and efficiency of fine-tuned open-source LLMs in real-world industrial applications. Our main contributions are summarized as follows:
  • Our study is one of the first (if not the first) to apply prompt engineering and fine-tuning techniques on open-source LLMs, exploring their potential for harmful data synthesis and improving toxic content detection.
  • We show that fine-tuned open-source LLMs deliver cost-effective and scalable solutions for automated content moderation, providing actionable insights for developing larger and more robust harmful content detection systems through synthetic data generation.
  • We also offer valuable insights into the practicalities of production deployment, highlighting real-world challenges and solutions.
@@ 2 Methodology
Our approach utilizes prompt engineering and fine-tuning techniques on open-source LLMs to enhance harmful data synthesis, thereby improving toxic content detection systems. The methodology comprises two key stages: prompt engineering and fine-tuning, as illustrated in Figure 1.
@@ 2.1 Prompt Engineering
where Role defines the context, Requirement specifies the task, and Few-Shot Examples provide examples to guide the model. In the initial phase of our study, we focused solely on prompt engineering. This involved creating various prompt templates tailored to elicit specific types of harmful data. Each prompt was carefully crafted following work by Zhou et al. (2023) to ensure clarity and relevance to the desired output. Despite the meticulous design of these prompts, we encountered significant challenges. The primary issue was the model’s inherent safety alignments, which are explicitly designed to prevent harmful content generation. While these alignments are crucial for ethical AI usage, they posed a hurdle for our specific goal of generating harmful data for augmentation purposes. Consequently, the generated data often lacked sufficient quality and diversity, hindering their utility in effective toxic content detection. To overcome these challenges, we experimented with various configurations, such as altering few-shot examples and adjusting requirement specificity. However, these adjustments yielded limited success in bypassing the internal safety mechanisms. After testing, the final prompt template are showed in Appendix A. Our findings showed that relying solely on prompt engineering was inadequate for our objectives. This realization prompted the integration of a second stage—fine-tuning—to enhance the model’s ability to generate high-quality harmful data by retraining its weights with carefully curated datasets.
@@ 2.2 Fine-Tuning
Based on the results from prompt engineering, we hypothesized that fine-tuning would yield a more effective solution. To achieve this, we employed LoRA methods Hu et al. (2022) for fine-tuning, updating the weights of $\mathcal{M}$ using the training dataset $\mathcal{D}_{\text{train}}$ to enhance the quality and diversity of $\mathcal{D}_{\text{harmful}}$. The fine-tuning objective, denoted as $\mathcal{L}$, minimizes the cross-entropy loss:
$\mathcal{L}=-\sum_{i=1}^{N}\sum_{j=1}^{|V|}y_{ij}\log(\hat{y}_{ij})$
where $y_{ij}$ represents the true label for the $j$-th vocabulary class of the $i$-th sample, $\hat{y}_{ij}$ is the predicted probability for the same, $|V|$ is the size of the vocabulary, and $N$ is the number of samples in the dataset.
After fine-tuning, Mistral demonstrated the best performance compared to the baselines; the detailed experimental results are presented in Section 4.
@@ 2.3 Algorithm
The overall algorithm for harmful data synthesis is outlined in Algorithm 1.
Algorithm 1 Harmful data synthesis using Open-Source LLMs
1:Open-source LLM $\mathcal{M}$, training dataset $\mathcal{D}_{\text{train}}$, prompt templates $\mathcal{P}$, downstream model $\mathcal{M}_{\text{down}}$
2:Augmented harmful data $\mathcal{D}_{\text{augmented}}$ and trained downstream model $\mathcal{M}_{\text{down}}$
@@
5:Fine-tune LLM $\mathcal{M}$ on $\mathcal{D}_{\text{train}}$ using LoRA method.
@@
7:     Generate harmful data samples $\mathcal{D}_{\text{harmful},i}$ using the fine-tuned LLM $\mathcal{M}$ and template $\mathcal{P}_{i}$
@@
9:Combine all generated samples: $\mathcal{D}_{\text{harmful}}=\bigcup_{i}\mathcal{D}_{\text{harmful},i}$
10:Combine $\mathcal{D}_{\text{harmful}}$ with $\mathcal{D}_{\text{train}}$ to form the augmented dataset: $\mathcal{D}_{\text{augmented}}=\mathcal{D}_{\text{train}}\cup\mathcal{D}_{\text{harmful}}$
11:for each batch $B_{j}$ in $\mathcal{D}_{\text{augmented}}$ do
12:     Train the downstream model $\mathcal{M}_{\text{down}}$ on $B_{j}$
@@
14:return $\mathcal{D}_{\text{augmented}}$, $\mathcal{M}_{\text{down}}$
Here, $\mathcal{D}_{\text{train}}$ represents the initial training dataset used for fine-tuning, $\mathcal{D}_{\text{harmful}}$ consists of unique harmful data samples generated by the fine-tuned LLM, and $\mathcal{D}_{\text{augmented}}$ is the final dataset combining $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{harmful}}$ for downstream model training. Duplicates in $\mathcal{D}_{\text{harmful}}$ are removed.
@@ 2.4 Evaluation
We evaluate the performance of the augmented data using F1-score and Accuracy metrics on a smaller MLP classifier (more detailed in Section 3.2).
@@ 3.1 Datasets
The following proprietary datasets are utilized for training and evaluating models designed to identify various types of harmful content. Each dataset consists of binary classification labels, categorizing texts as either positive (harmful) or negative (non-harmful). The labeling process involves multi-round human annotations to ensure accuracy and consistency across diverse content types. The datasets are sourced from online public datasets, company service collections, and other proprietary sources. Data samples are randomly selected from pool. Detailed data distributions are provided in Table 1. Additionally, a 2,500-entry evaluation dataset is prepared from each dataset for testing.
@@
Table 1: Distribution of Datasets
@@ 3.2 Models
n the first stage of our study, we employed prompt engineering to systematically evaluate the performance of six open-source LLMs: Mistral Jiang et al. (2023), LLaMA2 Touvron et al. (2023), Vicuna Chiang et al. (2023), Falcon Almazrouei et al. (2023), Bloom Workshop et al. (2023), and Gemma Team et al. (2024). Detailed results and analysis for Stage 1 are provided in Section 4.1. Based on the performance evaluation, Mistral and Vicuna were identified as the top-performing models in the prompt engineering stage, and were subsequently selected for the fine-tuning stage.
@@
In the second stage, we fine-tuned the Mistral and Vicuna models using the LoRA method Hu et al. (2022). After comparing latency and cost we decided to use a smaller model Mixtral for further exploration. Detailed results and analysis for Stage 2 are provided in Section 4.2.
@@ 3.3 Implementation
The fine-tuning process was conducted on an Azure cluster equipped with 4 * NVIDIA A100 GPUs, leveraging the LLaMA-Factory [^1] https://github.com/hiyouga/LLaMA-Factory framework Zheng et al. (2024). We employed a Supervised Fine-Tuning (SFT) methodology with a Low-Rank Adaptation (LoRA) setup Hu et al. (2022). Detailed training parameters, including batch size, learning rate, and optimizer settings, are provided in the Appendix B. The results of this fine-tuning process, along with performance metrics and qualitative analyses, are presented in Section 4.
@@
In our production environment, meeting strict inference time requirements is critical. Among the evaluated models, the Multilayer Perceptron (MLP) Gardner and Dorling (1998) stood out for its ability to fulfill these demands. For downstream model evaluation, in this paper we utilized a MLP classifier with two hidden layers, each comprising 600 neurons.
@@ 4 Result
This section presents the results of our study, showcasing the effectiveness of both prompt engineering and fine-tuning strategies in optimizing open-source LLMs for harmful data synthesis. Our evaluation employed a two-stage approach: Stage 1 focused on prompt engineering to assess the initial performance of various models, while Stage 2 involved fine-tuning the selected models to further enhance data quality and diversity.
@@ 4.1 Stage 1: Prompt Engineering
The first stage of our evaluation focused on assessing the performance of various Open Source LLMs for data synthesis using prompt engineering. The primary goal was to explore the quality and effectiveness of existing LLMs in generating augmented data for political content. Table 2 presents the performance comparison of the six different models evaluated in Stage 1. The success rates for political and hate content are reported, along with the results of human evaluation.
@@
Hate Mistral 0.608 0.706
GPT4(baseline) 0.628 0.721
@@
Sex_epoch5 0.755 0.788
Sex_epoch3 0.671 0.740
Sex_epoch2 0.653 0.731
Sex_epoch1 0.587 0.702
@@ 4.2 Stage 2: Fine-tuning
In the second stage, we began by fine-tuning the models exclusively on individual datasets and evaluating their performance on the same datasets. The results of this approach are detailed in Table 4. Inspired by Hui et al. (2024), Yu et al. (2024), and Qi et al. (2023), we subsequently adopted a data mixing strategy that combined multiple datasets for fine-tuning. The results of this multi-dataset fine-tuning are presented in Table 3. Additionally, we conducted an ablation study to examine the impact of different design choices, detailed in Appendix D. The "Mixture" versions represent fine-tuning on a mix of hate, violence, and sex datasets, while the "Mix_GPT" versions incorporate GPT-generated positive samples into the mixed data. For all fine-tuning experiments, we used 3000 positive and 3000 negative samples, following the Alpaca instruction template. Our findings reveal that fine-tuning significantly enhances the performance of LLMs in generating harmful data. The data mixing strategy helps in mitigating issues like data duplication and overfitting. The fine-tuned models exhibited higher F1 scores and accuracy across different datasets, demonstrating their robustness and effectiveness in real-world applications. One surprising observation was the notable improvement in model performance when using mixed data, suggesting that diverse training data enables better generalization and more reliable outputs. The Mixture models trained for Epoch 3 and Epoch 5 showed particularly strong results, emphasizing the importance of sufficient training epochs for optimal performance.